OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (3/2, -2) and (-3, 7/3)?

RhondaSommer (rhondasommer):

use the slope equeation. \[y2-y1\]----------\[x2-x1\]

OpenStudy (anonymous):

Ok thanks @RhondaSommer

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ \frac{3}{2}}}\quad ,&{\color{blue}{ -2}})\quad % (c,d) &({\color{red}{ -3}}\quad ,&{\color{blue}{ \frac{7}{3}}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (mathmale):

Nice graphics! jazzy-1: Can you now calculate the slope?